OpenStudy (anonymous):

Find the vertices and foci of the hyperbola with equation quantity x plus 5 squared divided by 36 minus the quantity of y plus 1 squared divided by 64 equals 1.

OpenStudy (anonymous):

@Saisuke<3 Do you know how to do this?

OpenStudy (anonymous):

@zepdrix @abb0t

zepdrix (zepdrix):

\[\Large\rm \frac{(x+5)^2}{6^2}-\frac{(y+1)^2}{8^2}=1\]Let's write it like this so we can easily identify our `a` and `b`. Hmm I don't remember hyperbolas so well :d gonna look it up real quick.

OpenStudy (anonymous):

yas thats how u set it up!!

OpenStudy (anonymous):

psh i forgot lol

OpenStudy (anonymous):

Ok. Thanks for being the one to finally get back to me @zepdrix

OpenStudy (anonymous):

#20minslater

zepdrix (zepdrix):

I'm trying to remember... we open in the direction of the positive one, ya? So since our x is the positive, we open left and right?

OpenStudy (anonymous):

I believe so

OpenStudy (anonymous):

It has been a while for me too @zepdrix I can't get through the whole problem by myself

zepdrix (zepdrix):

Our "center" is being shifted, so that will affect our vertex point here: